kyle plans to invest $8500, part at 4% simple interest and the rest at 5% simple interes what is the most that he can invest at 4% and still be guaranteed at least $405 in interest per yeay?
Answer by lwsshak3(11628)

kyle plans to invest $8500, part at 4% simple interest and the rest at 5% simple interes what is the most that he can invest at 4% and still be guaranteed at least $405 in interest per yeay?
**
let x= most of the $8500 than can be invested at 4%
8500-x=amt that will be invested at 5%
..
4%x+5%(8500-x)=405
.04x+425-.05x=405
.01x=20
x=2000
most of the $8500 than can be invested at 4%=$2000